# Fully Sparse 3D Occupancy Prediction

### v3

@@ Abstract
Occupancy prediction plays a pivotal role in autonomous driving. Previous methods typically construct dense 3D volumes, neglecting the inherent sparsity of the scene and suffering high computational costs. To bridge the gap, we introduce a novel fully sparse occupancy network, termed SparseOcc. SparseOcc initially reconstructs a sparse 3D representation from visual inputs and subsequently predicts semantic/instance occupancy from the 3D sparse representation by sparse queries. A mask-guided sparse sampling is designed to enable sparse queries to interact with 2D features in a fully sparse manner, thereby circumventing costly dense features or global attention. Additionally, we design a thoughtful ray-based evaluation metric, namely RayIoU, to solve the inconsistency penalty along depths raised in traditional voxel-level mIoU criteria. SparseOcc demonstrates its effectiveness by achieving a RayIoU of 34.0, while maintaining a real-time inference speed of 17.3 FPS, with 7 history frames inputs. By incorporating more preceding frames to 15, SparseOcc continuously improves its performance to 35.1 RayIoU without whistles and bells.
@@ 1 Introduction
Vision-centric 3D occupancy prediction [1] focuses on partitioning 3D scenes into structured grids from visual images. Each grid is assigned a label indicating if it is occupied or not. This task offers more geometric details than 3D detection and produces an alternative representation to LiDAR [61, 21, 58, 59, 60, 30, 31].
Existing methods [26, 14, 55, 44, 27] typically construct dense 3D features yet suffer from computational overhead (e.g. FPS on Tesla A100). However, dense representations are not necessary for occupancy predictions. Statistics in Fig. 1(a) reveal the geometry sparsity, that more than 90% of the voxels are empty. This manifests a large room in occupancy prediction acceleration by exploiting the sparsity. Some works [25, 17] explore the sparsity of 3D scenes, but they still rely on sparse-to-dense modules for dense predictions. This inspires us to seek a fully sparse occupancy network without any dense design.
In this paper, we propose SparseOcc, the first fully sparse occupancy network. As depicted in Fig. 1(b), SparseOcc includes two steps. First, it leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene in a coarse-to-fine manner. This only models non-free regions, saving computational costs significantly. Second, we design a mask transformer with sparse semantic/instance queries to predict masks and labels of segments from the sparse space. The mask transformer not only improves performance on semantic occupancy but also paves the way for panoptic occupancy. A mask-guided sparse sampling is designed to achieve sparse cross-attention in the mask transformer. As such, our SparseOcc fully exploits the sparse properties, with a fully sparse architecture free of any dense design like dense 3D features, sparse-to-dense modules, and global attention.
@@
Thanks to the sparsity design, SparseOcc achieves 34.0 RayIoU on Occ3D-nus [48], while maintaining a real-time inference speed of 17.3 FPS (Tesla A100, PyTorch fp32 backend), with 7 history frames inputs. By incorporating more preceding frames to 15, SparseOcc continuously improves its performance to 35.1 RayIoU, achieving state-of-the-art performance without whistles and bells. The comparison between SparseOcc with previous methods in terms of performance and efficiency is shown in Fig. 1(c).
@@
We propose SparseOcc, the first fully sparse occupancy network without any time-consuming dense designs. It achieves 34.0 RayIoU on Occ3D-nus benchmark with an inference speed of 17.3 FPS.
@@ 2 Related Work
Occupancy Networks were originally proposed by Mescheder et al. [36, 41], with a focus on continuous object representations in 3D space. Recent variations in occupancy networks [1, 4, 44, 48, 52, 54, 56] shift their attention towards reconstructing 3D space [23, 22, 26, 14, 13, 16, 15, 53, 33, 34, 32, 51, 57] and predicts voxel-level semantic information from image inputs. MonoScene [4] achieves scene occupancy completion through a 2D and a 3D UNet [42] connected by a sight projection module. OccNet [44] applies universal occupancy features to various downstream tasks and introduces the OpenOcc benchmark. SurroundOcc [55] proposes a coarse-to-fine architecture. However, the computational burden of handling a large number of voxel queries remains a challenge. TPVFormer [17] proposes using tri-perspective view representations to supplement vertical structural information, but this inevitably leads to information loss. VoxFormer [25] initializes sparse queries based on monocular depth prediction. However, VoxFormer is not fully sparse as it still requires a sparse-to-dense MAE [11] module to complete the scene. Some methods emerged in the CVPR 2023 occupancy challenge [27, 39, 9], but none of them exploits a fully sparse design. In this paper, we make the first step to explore the fully sparse architecture for 3D occupancy predictions from camera inputs.
@@
As a related task to 3D occupancy prediction, 3D reconstruction recovers the 3D geometry from multiple posed images. Recent methods focus on more compact and efficient end-to-end 3D reconstruction pipelines [38, 46, 2, 45, 10]. Atlas [38] extracts features from multi-view input images and maps them to 3D space to construct the truncated signed distance function [8]. NeuralRecon [46] directly reconstructs local surfaces as sparse TSDF volumes and uses a GRU-based TSDF fusion module to fuse features from previous fragments. VoRTX [45] utilizes transformers to address occlusion issues in multi-view images. CVRecon [10] takes a different approach by avoiding occlusion-sensitive perspective mappings of 2D to 3D and directly uses cost volumes to establish 3D features from 2D image features.
@@
Recently, unified segmentation models have been widely studied to handle semantic and instance segmentation concurrently. Cheng et al. first proposes MaskFormer [7] for unified segmentation in terms of model architecture, loss functions, and training strategies. Mask2Former [6] then introduces masked attention, with restricted receptive fields on instance masks, for better performance. Later on, Mask3D [43] successfully extends the mask transformer for point cloud segmentation with state-of-the-art performance. OpenMask3D [47] further achieves the open-vocabulary 3D instance segmentation task and proposes a model for zero-shot 3D segmentation.
@@ 3.1 Sparse Voxel Decoder
Since 3D occupancy ground truths [48, 44, 55, 52] are dense 3D volume with shape as (e.g. 20020016), existing methods typically build a dense 3D feature of shape , but suffer from computational overhead. In this paper, we argue that such dense representation is not necessary for occupancy prediction. As in our statistics, we find that over 90% of the voxels in the scene are free. This motivates us to explore a sparse 3D representation that only models the non-free areas of the scene, thereby saving computational resources.
@@
Our designed sparse voxel decoder is shown in Fig. 3. In general, it follows a coarse-to-fine structure but only models the non-free regions. The decoder starts from a set of coarse voxel queries equally distributed in the 3D space (e.g. 2525). In each layer, we first upsample each voxel by 2, e.g. a voxel with size will be upsampled into 8 voxels with size . Next, we estimate an occupancy score for each voxel and conduct pruning to remove useless voxel grids. Here we have two approaches for pruning: one is based on a threshold (e.g., only keeps score 0.5); the other is by top- selection. In our implementation, we simply keep voxels with top- occupancy scores for training efficiency. is a dataset-related parameter, obtained by counting the maximum number of non-free voxels in each sample at different resolutions. The voxel tokens after pruning will serve as the input for the next layer.
@@
Within each layer, we use a transformer-like architecture to handle voxel queries. The concrete architecture is inspired by SparseBEV [32], a detection method using a sparse scheme. To be specific, in layer with voxel queries described by 3D locations and a -dim content vector, we first use self-attention to aggregate local and global features for those query voxels. Then, a linear layer is used to generate 3D sampling offsets for each voxel query from the associated content vector. These sampling offsets are utilized to transform voxel queries to obtain reference points in global coordinates. We finally project those sampled reference points to multi-view image space for integrating image features by adaptive mixing [49, 18].
@@
Previous dense occupancy methods [26, 44, 14] typically warp the history BEV/3D feature to the current timestamp, and use deformable attention [62] or 3D convolutions to fuse temporal information. However, this approach is not directly applicable in our case due to the sparse nature of our 3D features. To handle this, we leverage the flexibility of the aforementioned global sampled reference points by warping them to previous timestamps to sample history multi-view image features. The sampled multi-frame features are then stacked and aggregated by adaptive mixing so as for temporal modeling.
@@
We compute loss for sparse voxels from each layer. We use binary cross entropy (BCE) loss as the supervision, given that we are reconstructing a class-agnostic sparse occupancy space. Only the kept sparse voxels are supervised, while the discarded regions during pruning in earlier stages are ignored.
@@ 3.2 Mask Transformer
Our mask transformer is inspired by Mask2Former [6], which uses sparse semantic/instance queries decoupled by binary mask query and content vector . The mask transformer consists of three steps: multi-head self attention (MHSA), mask-guided sparse sampling, and adaptive mixing. MHSA is used for the interaction between different queries as the common practice. Mask-guided sparse sampling and adaptive mixing are responsible for the interaction between queries and 2D image features.
@@
For class prediction, we apply a linear classifier with a sigmoid activation based on the query embeddings . For mask prediction, the query embeddings are converted to mask embeddings by a MLP. The mask embeddings have the same shape as query embeddings and are dot-producted with the sparse voxel embeddings to produce mask predictions. Thus, the prediction space of our mask transformer is constrained to the sparsified 3D space from the sparse voxel decoder, rather than the full 3D scene. The mask predictions will serve as the mask query for the next layer.
@@
Follow MaskFormer [7], we match the ground truth with the predictions using Hungarian matching. Focal loss is used for classification, while a combination of DICE loss [37] and BCE mask loss is used for mask prediction. Thus, the total loss of SparseOcc is composed of:
@@ 4.1 Revisiting the Voxel-level mIoU
The Occ3D dataset [48], along with its proposed evaluation metrics, are widely recognized as benchmarks in this field. The ground truth occupancy is reconstructed from LiDAR measurements, and the mean Intersection over Union (mIoU) at the voxel level is employed to assess performance.
However, due to factors such as distance and occlusion, accumulated LiDAR point clouds are imperfect. Some areas unscanned by LiDAR are marked as free, resulting in fragmented instances, as shown in Fig. 4(a). This raises problems of label inconsistency. Previous efforts in solving the evaluation issue, e.g. Occ3D, use a binary visible mask that indicates whether the voxels are observed in the current camera view. However, we found that only calculating mIoU on the observed voxel position could still cause ambiguity and be easily hacked. As illustrated in Fig. 4, RenderOcc [39] generated a cracked but thick surface, approximately unusable for downstream task. Nevertheless, RenderOcc’s masked mIoU is much higher than BEVFormer’s prediction [26], which is more stable and clean.
The misalignment between qualitative and quantitative results is caused by the inconsistency along the depth direction of the visible mask. As shown in Fig. 5, this toy example reveals several issues with the current evaluation metrics:
@@
If the model fills all areas behind the surface, it inconsistently penalizes the prediction of depth. The model can obtain a higher IoU by filling all areas behind the surface and predicting a closer depth. This thick surface issue is very common in models that use the visible mask or 2D supervision.
@@
If the predicted occupancy is a thin surface, the penalty is overly strict, as a deviation of one voxel will lead to an IoU of zero.
@@
The visible mask only considers the visible area at the current moment, thereby reducing Occupancy to a depth estimation task with categories and overlooking the crucial ability to complete scenes beyond the visible region.
@@ 4.2 Mean IoU by Ray Casting
To address the above issues, we propose a new evaluation metric: Ray-level mIoU (RayIoU for short). In RayIoU, the elements of the set are now query rays, not voxels. We emulate LiDAR by projecting query rays into the predicted 3D occupancy volume. For each query ray, we compute the distance it travels before intersecting any surface and retrieve the corresponding class label. We then apply the same procedure to the ground-truth occupancy to obtain the ground-truth depth and class label. In case a ray doesn’t intersect with any voxel present in the ground truth, it will be excluded from the evaluation process.
As shown in Fig. 6(a), the original LiDAR rays in a real dataset tend to be unbalanced from near to far. Thus, we resample the LiDAR rays to balance the distribution on different distance, as shown in Fig. 6(b). For the near field, we modify the LiDAR ray channels to achieve equal-distant spacing when projected onto the ground plane. In the far field, we increase the angular resolution of the ray channels to ensure a more uniform data density across varying ranges. Moreover, our query ray can originate from the LiDAR position at the current, past, or future moments of the ego path. As shown in Fig. 6, temporal casting allows us to better evaluate the scene completion performance while ensuring that the task remains well-posed.
A query ray is classified as a true positive (TP) if the class labels coincide and the L1 error between the ground-truth depth and the predicted depth is less than a certain threshold (e.g., 2m). Let be he number of classes,
where , , and correspond to the number of true positive, false positive, and false negative predictions for class .
@@
Since the query ray only calculates the distance it touches the first voxel, the model cannot obtain a higher IoU by filling the area behind the surface.
@@
RayIoU determines TP through the distance threshold, mitigating the overly harsh nature of voxel-level mIoU.
@@
The query ray can originate from any position in the scene, thereby considering the model’s completion ability and preventing the reduction of occupancy to depth estimation.

### v4

@@ Abstract
Occupancy prediction plays a pivotal role in autonomous driving. Previous methods typically construct dense 3D volumes, neglecting the inherent sparsity of the scene and suffering from high computational costs. To bridge the gap, we introduce a novel fully sparse occupancy network, termed SparseOcc. SparseOcc initially reconstructs a sparse 3D representation from camera-only inputs and subsequently predicts semantic/instance occupancy from the 3D sparse representation by sparse queries. A mask-guided sparse sampling is designed to enable sparse queries to interact with 2D features in a fully sparse manner, thereby circumventing costly dense features or global attention. Additionally, we design a thoughtful ray-based evaluation metric, namely RayIoU, to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria. SparseOcc demonstrates its effectiveness by achieving a RayIoU of 34.0, while maintaining a real-time inference speed of 17.3 FPS, with 7 history frames inputs. By incorporating more preceding frames to 15, SparseOcc continuously improves its performance to 35.1 RayIoU without bells and whistles.
@@ 1 Introduction
Vision-centric 3D occupancy prediction [1] focuses on partitioning 3D scenes into structured grids from visual images. Each grid is assigned a label indicating if it is occupied or not. This task offers more geometric details than 3D object detection and produces an alternative representation to LiDAR-based perception [61, 22, 58, 59, 60, 29, 30].
Existing methods [26, 15, 55, 43, 27] typically construct dense 3D features yet suffer from computational overhead (e.g., FPS on the Tesla A100 GPU). However, dense representations are not necessary for occupancy prediction. We statistic the geometry sparsity and find that more than 90% of the voxels are empty. This manifests a large room in occupancy prediction acceleration by exploiting the sparsity. Some works [25, 18] explore the sparsity of 3D scenes, but they still rely on sparse-to-dense modules for dense predictions. This inspires us to seek a pure sparse occupancy network without any dense design.
In this paper, we propose SparseOcc, the first fully sparse occupancy network. As depicted in Fig. 1 (a), SparseOcc includes two steps. First, it leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene in a coarse-to-fine manner. This only models non-free regions, saving computational costs significantly. Second, we design a mask transformer with sparse semantic/instance queries to predict masks and labels of segments from the sparse space. The mask transformer not only improves performance on semantic occupancy but also paves the way for panoptic occupancy. A mask-guided sparse sampling is designed to achieve sparse cross-attention in the mask transformer. As such, our SparseOcc fully exploits the sparse property and gets rid of any dense design like dense 3D features, sparse-to-dense modules, and global attention.
@@
Thanks to the sparsity design, SparseOcc achieves 34.0 RayIoU on Occ3D-nuScenes [48], while maintaining a real-time inference speed of 17.3 FPS (Tesla A100, PyTorch fp32 backend), with 7 history frames inputs. By incorporating more preceding frames to 15, SparseOcc continuously improves its performance to 35.1 RayIoU, achieving state-of-the-art performance without bells and whistles. The comparison between SparseOcc with previous methods in terms of performance and efficiency is shown in Fig. 1 (b).
@@
We propose SparseOcc, the first fully sparse occupancy network without any time-consuming dense designs. It achieves 34.0 RayIoU on Occ3D-nuScenes benchmark with an real-time inference speed of 17.3 FPS.
@@ 2 Related Work
The occupancy network is originally proposed by Mescheder et al. [35, 40], focusing on continuous object representations in 3D space. Recent variations [1, 4, 43, 48, 52, 54, 11, 56] mostly draw inspiration from Bird’s Eye View (BEV) perception [24, 23, 26, 15, 14, 17, 16, 53, 32, 33, 31, 51, 57] and predicts voxel-level semantic information from image inputs. For instance, MonoScene [4] estimates occupancy through a 2D and a 3D UNet [41] connected by a sight projection module. SurroundOcc [55] proposes a coarse-to-fine architecture. However, the large number of voxel queries is computationally heavy. TPVFormer [18] proposes tri-perspective view representations to supplement vertical structural information, but this inevitably leads to information loss. VoxFormer [25] initializes sparse queries based on monocular depth prediction. Nevertheless, VoxFormer is not fully sparse as it still requires a sparse-to-dense MAE [12] module to complete the scene. Some methods emerged in the CVPR 2023 occupancy challenge [27, 38, 9], but none of them exploits a fully sparse design. In this paper, we make the first step to explore the fully sparse architecture for 3D occupancy prediction from camera-only inputs.
Sparse Architectures for 3D Vision.
Sparse architectures find widespread adoption in LiDAR-based reconstruction [46] and perception [7, 61, 58, 59], leveraging the inherent sparsity of point clouds. However, when it comes to vision-to-3D tasks, a direct adaptation is not feasible due to the absence of point cloud inputs. A prior work, SparseBEV [31], proposes a fully sparse architecture for camera-based 3D object detection. Nevertheless, directly adapting this approach is non-trivial because 3D object detection focuses on a sparse set of objects, whereas 3D occupancy requires dense predictions for each voxel. Consequently, designing a fully sparse architecture for 3D occupancy prediction remains a challenging task.
@@
As a related task to 3D occupancy prediction, 3D reconstruction recovers the 3D geometry from multiple posed images. Recent methods focus on more compact and efficient end-to-end 3D reconstruction pipelines [37, 45, 2, 44, 10]. Atlas [37] extracts features from multi-view input images and maps them to 3D space to construct the truncated signed distance function [8]. NeuralRecon [45] directly reconstructs local surfaces as sparse TSDF volumes and uses a GRU-based TSDF fusion module to fuse features from previous fragments. VoRTX [44] utilizes transformers to address occlusion issues in multi-view images.
@@
Recently, unified segmentation models have been widely studied to handle semantic and instance segmentation concurrently. Cheng et al. first propose MaskFormer [6] for unified segmentation in terms of model architecture, loss functions, and training strategies. Mask2Former [5] then introduces masked attention, with restricted receptive fields on instance masks, for better performance. Later on, Mask3D [42] successfully extends the mask transformer for point cloud segmentation with state-of-the-art performance. OpenMask3D [47] further achieves the open-vocabulary 3D instance segmentation task and proposes a model for zero-shot 3D segmentation.
@@ 3.1 Sparse Voxel Decoder
Since 3D occupancy ground truth [48, 43, 55, 52] is a dense volume with dimensions (e.g., 20020016), existing methods typically build a dense 3D feature of shape , but suffer from computational overhead. In this paper, we argue that such dense representation is not necessary for occupancy prediction. As in our statistics, we find that over 90% of the voxels in the scene are free. This motivates us to explore a sparse 3D representation that only models the non-free areas of the scene, thereby saving computational resources.
@@
Our designed sparse voxel decoder is shown in Fig. 3. In general, it follows a coarse-to-fine structure but only models the non-free regions. The decoder starts from a set of coarse voxel queries equally distributed in the 3D space (e.g., 2525). In each layer, we first upsample each voxel by 2, e.g., a voxel with size will be upsampled into 8 voxels with size . Next, we estimate an occupancy score for each voxel and conduct pruning to remove useless voxel grids. Here we have two approaches for pruning: one is based on a threshold (e.g., only keeps score 0.5); the other is by top- selection. In our implementation, we simply keep voxels with top- occupancy scores for training efficiency. is a dataset-related parameter, obtained by counting the maximum number of non-free voxels in each sample at different resolutions. The voxel tokens after pruning will serve as the input for the next layer.
@@
Within each layer, we use a transformer-like [50] architecture to handle voxel queries. The concrete architecture is inspired by SparseBEV [31], a detection method using a sparse scheme. To be specific, in layer with voxel queries described by 3D locations and a -dim content vector, we first use self-attention to aggregate local and global features for those query voxels. Then, a linear layer is used to generate 3D sampling offsets for each voxel query from the associated content vector. These sampling offsets are utilized to transform voxel queries to obtain reference points in global coordinates. We finally project those sampled reference points to multi-view image space for integrating image features by adaptive mixing [49, 19]. In summary, our approach differs from SparseBEV by shifting the query formulation from pillars to 3D voxels. Other components such as self attention, adaptive sampling and mixing are directly borrowed.
@@
Previous dense occupancy methods [26, 15] typically warp the history BEV/3D feature to the current timestamp, and use deformable attention [62] or 3D convolutions to fuse temporal information. However, this approach is not directly applicable in our case due to the sparse nature of our 3D features. To handle this, we leverage the flexibility of the aforementioned global sampled reference points by warping them to previous timestamps to sample history multi-view image features. The sampled multi-frame features are stacked and aggregated by adaptive mixing so as for temporal modeling.
@@
We compute loss for the sparsified voxels from each layer. We use binary cross entropy (BCE) loss as the supervision, given that we are reconstructing a class-agnostic sparse occupancy space. Only the kept sparse voxels are supervised, while the discarded regions during pruning in earlier stages are ignored.
@@ 3.2 Mask Transformer
Our mask transformer is inspired by Mask2Former [5], which uses sparse semantic/instance queries decoupled by binary mask queries and content vectors . The mask transformer consists of three steps: multi-head self attention (MHSA), mask-guided sparse sampling, and adaptive mixing. MHSA is used for the interaction between different queries as the common practice. Mask-guided sparse sampling and adaptive mixing are responsible for the interaction between queries and 2D image features.
@@
For class prediction, we apply a linear classifier with a sigmoid activation based on the query embeddings . For mask prediction, the query embeddings are converted to mask embeddings by an MLP. The mask embeddings have the same shape as query embeddings and are dot-producted with the sparse voxel embeddings to produce mask predictions. Thus, the prediction space of our mask transformer is constrained to the sparsified 3D space from the sparse voxel decoder, rather than the full 3D scene. The mask predictions will serve as the mask queries for the next transformer layer.
@@
Following MaskFormer [6], we match the ground truth with the predictions using Hungarian matching. Focal loss is used for classification, while a combination of DICE loss [36] and BCE mask loss is used for mask prediction. Thus, the total loss of SparseOcc is composed of four parts:
@@ 4.1 Revisiting the Voxel-level mIoU
The Occ3D dataset [48], along with its proposed evaluation metrics, are widely recognized as benchmarks in this field. The ground truth occupancy is reconstructed from LiDAR point clouds, and the mean Intersection over Union (mIoU) at the voxel level is employed to assess performance. Due to factors such as distance and occlusion, the accumulated point clouds are not perfect. Some areas unscanned by LiDAR are marked as free, resulting in fragmented instances. This raises the problem of label inconsistency. To solve this problem, Occ3D uses a binary visible mask that indicates whether a voxel is observed in the current camera view. Only the observed voxels contribute to evaluation.
However, we found that solely calculating mIoU on the observed voxel positions remains vulnerable and can be hacked by predicting a thicker surface. Dense methods (e.g., BEVFormer [26]) can easily achieve this by training with the visible mask. During training, the area behind the surface lacks supervision, causing the model to fill it with duplicated predictions, resulting in a thicker surface. As an example, consider BEVFormer, which generates a thick and noisy surface when trained with the visible mask (see Fig. 4). Despite this, its performance exhibits an unreasonably inflated improvement (+515 mIoU) under the current evaluation protocol.
The misalignment between qualitative and quantitative results is caused by the inconsistent penalty along the depth direction. A toy example in Fig. 5 reveals several issues with the current metrics:
@@
If the model fills all areas behind the surface, it inconsistently penalizes depth predictions. The model can obtain a higher IoU by filling all areas behind the surface and predicting a closer depth. This thick surface issue is very common in dense models trained with visible masks or 2D supervision.
@@
If the predicted occupancy represents a thin surface, the penalty becomes overly strict. Even a deviation of just one voxel results in an IoU of zero.
@@
The visible mask only considers the visible area at the current moment, reducing occupancy prediction to a depth estimation task and overlooking the scene completion ability.
@@ 4.2 Mean IoU by Ray Casting
To address the above issues, we propose a new evaluation metric: Ray-level mIoU (RayIoU for short). In RayIoU, the set elements are query rays rather than voxels. We emulate LiDAR behavior by projecting query rays into the predicted 3D occupancy volume. For each query ray, we compute the distance it travels before intersecting any surface and retrieve the corresponding class label. We then apply the same procedure to the ground-truth occupancy to obtain the ground-truth depth and class label. In case a ray does not intersect with any voxel present in the ground truth, it will be excluded from the evaluation process.
As shown in Fig. 6 (a), the raw LiDAR rays in a real dataset tend to be unbalanced from near to far. Thus, we resample the rays to achieve a balanced distribution across different distances (Fig. 6 (b)). In the near field, we modify the ray channels to achieve equal-distant spacing when projected onto the ground plane. In the far field, we increase the angular resolution of the ray channels to ensure a more uniform data density across varying ranges. Moreover, our query ray can originate from the LiDAR position at the current, past, or future moments of the ego path. Temporal casting (Fig. 6 (c)) allows us to evaluate scene completion performance while maintaining a well-posed task.
A query ray is classified as a true positive (TP) if the class labels coincide and the L1 error between the ground-truth depth and the predicted depth is less than a certain threshold (e.g., 2m). Let be the number of classes, then RayIoU is calculated as follows:
where , and correspond to the number of true positive, false positive, and false negative predictions for class .
@@
Since the query ray calculates the distance to the first voxel it touches, the model cannot obtain a higher IoU by predicting a thicker surface.
@@
RayIoU determines true positives based on a distance threshold, which mitigates the overly strict nature of voxel-level mIoU.
@@
The query ray can originate from any position in the scene. This flexibility allows RayIoU to consider the model’s scene completion ability, preventing the reduction of occupancy estimation to mere depth prediction.